The rotating loop in an AC generator is a square 15.0 cm on a side. It is rotated at 70.0 Hz in a uniform field of 0.800 Calculate the following quantities as functions of time t, where t is in seconds. (a) the flux through the loop mT.m (b) the emf induced in the loop (c) the current induced in the loop for a loop resistance of 1.50 n (d) the power delivered to the loop W (e) the torque that must be exerted to rotate the loop mN m
Answer #1

a)

phi = B a^2 = 0.8* 0.15^2 =18 m T/m^2

======

b)

emf induced

E = phi * w = 0.018* (2* pi * 70) = 7.91 V

======

c)

i = e/R = 5.2752 A

=====

d)

P = e i = 41.73 W

=====

e)

T = B i A = 95 mN-m

=====

comment before rate in case any doubt, will reply for sure.. goodluck